OpenStudy (anonymous):

If we list all the natural numbers below 10 that are multiples of 3 or 5, we get 3, 5, 6 and 9. The sum of these multiples is 23. Find the sum of all the multiples of 3 or 5 below 1000.

OpenStudy (zzr0ck3r):

what class are you doing this in?

OpenStudy (anonymous):

See, find the sum of all the multiples of 3 first that are coming under 1000: That is: 3 + 6 + 9+............................. +999 This is An Arithmetic Progression having a = 3 and d = 3 Last term is : 999 = a + (n - 1)d 999 = 3 + 3n - 3 3n = 999 n = 333 So you can find the sum by using : \[\huge Sum = \frac{n}{2}(a + l)\] Where l = Last term = 999 So, \[Sum = \frac{333}{2}(3 + 999) = 333 \times 501 = 166833\] Now doing the same procedure for: 5 + 10 + 15+ ........................ + 995 a = 5, d= 5 l= 995 So, \[Sum = \frac{199}{2}(5 + 995) = 199 \times 500 = 99500\] Adding the two sums: 166833 + 99500 = 266333

OpenStudy (anonymous):

There is one another method which is quite simpler then this.. If you want to know then tell I will explain it to you...

OpenStudy (anonymous):

thnx for the reply! and yes @waterineyes i would also like to know it

OpenStudy (anonymous):

See, you have to find first for: 3 + 6 + 9 + 12+ .................. + 999 Take 3 common: 3(1 + 2 + 3 + 4 + ......+333) It is sum of first 333 natural numbers: Its formula for sum is given by: \[Sum = \frac{n(n+1)}{2}\] Here n= 333 put and tell me what will be the sum..

OpenStudy (anonymous):

55611

OpenStudy (anonymous):

So, this is value of brackets; so total sum = 3(55611) = ???

OpenStudy (anonymous):

166833

OpenStudy (anonymous):

Can you see by first method we have also the same result??

OpenStudy (anonymous):

yes thnx very much!

OpenStudy (anonymous):

Now solve for 5 + 10 + 15 + 20 ................+ 995 take 5 common and do the same steps as we have just done in finding the sum of 3 multiples..

OpenStudy (anonymous):

ok thnx very much!

OpenStudy (anonymous):

Welcome dear..